# RAPTOR: Practical Numerical Profiling of Scientific Applications

### v1

@@ 6.3. Mem-mode debugging
In this section, we demonstrate how the debugging output generated by RAPTOR in mem-mode can identify sections of code that may be unstable after truncatation. A domain scientist can use this information to understand which parts of the code contribute most to the overall accuracies when truncated.
To recap Section 3.5: mem-mode uses shadow variables that track an reference result for every operation as if the entire application had been run in FP64 up to that point. This feature can be used to pinpoint exactly where calculations start to deviate from the reference (above a threshold). RAPTOR flags these occurrences of inexact operation results, groups them by (debug) location in the code, and dumps the collected statistics when instructed by the user. In practice, certain instances of operations get flagged more often than others and flagged operations tend to be close together in the code. Thus, the output of mem-mode builds a heatmap of code locations that do not react well to truncation.
We apply mem-mode debugging to the Sedov problem with a newer, more modular hydrodynamics solver, Spark (Couch et al., 2021). The modularity of Spark makes it easier to ascribe a certain task to each part of the code. For example, two important components of the solver are the reconstruction algorithm and the Riemann solver. The reconstruction algorithm approximates the variation of the solution within each cell with a profile instead of using a constant value. The Riemann solver handles discontinuous solutions in shocks. For this experiment, we keep the timestep of the solver constant to ensure that the dynamic time-stepping algorithm does not interfere with our experiment by compensating for inaccuracies resulting from truncation.
@@
After we truncate the hydrodynamics, RAPTOR flags a number of operations in the reconstruction algorithm. This can be interpreted in two ways. Either, (a) the reconstruction algorithm does not respond well to truncation and should therefore be run at full precision, or (b) another part of the code is responsible for introducing small floating-point errors that only get flagged when they are amplified in the reconstruction algorithm, for example due to multiplication or cancellation. If we exclude reconstruction from truncation (but still truncate the remaining hydrodynamics code), then the errors for both density and x-velocity decrease by a small amount. Now, a number of new operations get flagged in the Riemann solver and a single operation gets flagged in the function responsible for updating the solution of the simulation. Adding the Riemann solver to the list of excluded functions drastically worsens errors, while adding the update function leaves the errors essentially unchanged.
These results indicate that no specific part of the Spark solver is responsible for numerical sensitivity in the Sedov workload and it is difficult to decide in advance whether a code section should be truncated or not. Instead, an effective truncation strategy will dynamically turn on and off based on the local smoothness of the solution in the physical regime. This small example reaffirms the need for a tool, such as RAPTOR, that can help the users to understand and reason about the parts of a code that may contribute the most to floating-point errors.
@@ 7. Discussion and Outlook
We used RAPTOR to test our hypotheses about how to truncate a number of non-linear multi-physics workloads while keeping errors in check. For Sedov, we found that restricting truncation to coarse blocks worked well to keep errors low. Meanwhile for Sod, the same strategy resulted in much higher errors. These two results match Hypothesis 1. However, our Hypothesis 2 was falsified when we tried to truncate the EOS module in Cellular and found that it only converges when we run it with double precision. Finally, we evaluated the effect of precision on the evolution of the Bubble workload and found that position and shape of the split bubbles depends on the strategy and level of truncation, as the scientific intuition anticipated.
@@ 7.1. Overhead
RAPTOR’s overhead for typical executions is shown in Table 3. The overhead roughly correlates to the proportion of operations we truncate, and in the simplest op-mode it can get up to 36$\times$ slower (with lower overheads for lower proportions). Since our tool is not meant to be used in production runs, and is only meant for experimentation and reasoning about precision, we believe that the overhead is within acceptable ranges that make it useful.
Figure 8. Estimated speedup of Sod (cf. §4.2 and Figure 7(b)) for different truncation strategies, according to our hardware model under a compute-bound or memory-bound assumption. Here, our model predicts a compute-bound scenario.
@@ 7.2. Use in Hardware Co-design
We want to also highlight other adjunct usecases for RAPTOR, e.g., as part of the hardware co-design (w.r.t. floating-point capabilities).
@@
Assume, we have a simple model for a hypothetical processor with a fixed portion of chip area dedicated to floating-point processing. Because the workloads, which we investigated earlier, are all double precision floating-point, and we truncate a portion of them to a lower precision, we can assume that our CPU contains FPUs for double precision and one lower precision. For simplicity, we will assume that the areas dedicated to each unit remain the same.
Table 4. Performance density of FPUs for various precisions (data from FPNew (Mach et al., 2020))
@@
We estimate how much processing power in floating-point operations per second (FLOP/s) a unit area of chip footprint provides for a given precision [^24] For simplicity, we assume that floating point units with different precision are completely independent. In reality, floating point units for different precisions often share some components.. We get an estimate for the performance density of an FPU using data from an open-source RISCV FPU implementation (i.e., FPNew (Mach et al., 2020)), which is summarized in Table 4. We extrapolate these values to get a performance density estimate for FPUs of any given precision.
Our theoretical CPU has a separate FPU for the two supported precisions (called $M_{\mathrm{dbl}}$ and $M_{\mathrm{low}}$), which, respectively, occupy an area of $A_{\mathrm{dbl}}$ and $A_{\mathrm{low}}$ and have a performance density of $P_{\mathrm{dbl}}$ and $P_{\mathrm{low}}$ operations per second per unit area (obtained above using extrapolation). Therefore, the processing capability for a specific precision is bound by a peak of $A_{i}\times P_{i}$ operations per unit time. To estimate how much area each of the two FPUs will occupy, we assume a typical CPU configuration where we have double and single precision capabilities with a ratio of $1:2$ (e.g. Fugaku’s A64FX (fug, [n. d.])). Using the estimated performance densities $P_{i}$, and the ratio of the compute power, we get the ratio for the areas $A_{\mathrm{dbl}}:A_{\mathrm{low}}=1.39$.
@@
To get a prediction for whether a computation is compute- or memory-bound, we build a roofline model of our processor, assuming a bandwidth of $1024$ GB/s (e.g., Fugaku’s (fug, [n. d.])).
@@ Results
Using the above model and the data collected from our experiments described earlier, we obtain estimated speedups for different configurations of Sod and present the results in Figure 8. Due to this being a computation-heavy application that has a high operation intensity (operations per byte), our roofline model predicted a compute-bound scenario. The irregularities around the small mantissas are due to the AMR behavior changing and resulting in a different ratios of truncated and non-truncated operations. With insight from domain scientists on acceptable levels of error obtained from Figure 7(b), this information can be used to obtain an optimal FPU configuration for hardware co-design. Given that RAPTOR simplifies the collection of such information, collaborating with scientists for gathering data on the numerical behavior of software can become a powerful way to enable supercomputing centers to make informed decisions about future procurements appropriate for their workloads.
@@ 7.3. Current Limitations
For our proof-of-concept, we omitted a few quality-of-life features. For example, not all elementary functions are implemented, but adding additional functions is trivial if MPFR already supports them, otherwise minor engineering is required. RAPTOR does not truncate function calls to external, pre-compiled libraries [^25] This can be mitigated by making the function definitions available, either by providing source code or static libraries compiled with link-time optimization (LTO). or calls to dynamic function pointers. We have not tested OpenMP support or thread-safety in general, original floating-point operands have to be FP32 or FP64, and we do not support handwritten vector assembly. These limitations can be overcome with small engineering effort. As shown in Figure 4, the user has to manually change a few lines of code. To improve the user experience, RAPTOR can be enhanced for easier region annotation (e.g. #pragma trunc...) and to support function filtering using a configuration file (similar to profilers). Programming languages other than C/C++ and Fortran could be supported if they lower to LLVM IR. Lastly, our truncation parameters are compile-time constants, but deciding the truncation level at runtime can be achieved by compiling multiple function pointers for different truncations and conditionally using them.
@@ 8. Conclusion
This paper introduces a novel numerical profiling approach and accompanying tool, called RAPTOR, which overcomes the limitations of state-of-the-art tools in this category. Using an LLVM compiler pass and MPFR, we are able to transparently change the floating-point precision by modifying the exponent and/or mantissa of selected code regions. We create hypotheses for multiple, different workloads in the Flash-X framework for multi-physics simulations. Using our RAPTOR tool, we demonstrate how we are able to analyze and confirm/falsify theses hypotheses based on collected data for the non-linear solvers which are usually hard to reason about. Our hypotheses for Sod, Sedov and Bubble experiments prove to be correct. However, our intuition turns out to be incorrect for the Cellular experiment. Currently, we can apply our tool to C/C++ and Fortran codes, and hence it supports the majority of (large-scale) scientific workloads run on supercomputers. While we outlined a few limitations, none of them are fundamental to our approach, and, as we mention, these can be mitigated with additional engineering effort. RAPTOR is a step forward towards reasoning about mixed/low precision in our simulations, and we hope that it finds wide adoption in the scientific-computing community. We make our RAPTOR tool accessible at https://github.com/RAPTOR-FP/RAPTOR for others to use.

### v2

@@ 6.3. Mem-mode debugging
In this section, we demonstrate how the debugging output generated by RAPTOR in mem-mode can identify sections of code that may be unstable after truncation. A domain scientist can use this information to understand which parts of the code contribute most to the overall inaccuracies when truncated.
To recap Section 3.5: mem-mode uses shadow variables that track a reference result for every operation as if the entire application had been run in FP64 up to that point. This feature can be used to pinpoint exactly where calculations start to deviate from the reference (above a threshold). RAPTOR flags these occurrences of inexact operation results, groups them by (debug) location in the code, and dumps the collected statistics when instructed by the user. In practice, certain instances of operations get flagged more often than others and flagged operations tend to be close together in the code. Thus, the output of mem-mode builds a heatmap of code locations that do not react well to truncation.
We apply mem-mode debugging to the Sedov problem with a newer, more modular hydrodynamics solver, Spark (Couch et al., 2021). The modularity of Spark makes it easier to ascribe a certain task to each part of the code. For example, two important components of the solver are the reconstruction algorithm and the Riemann solver. The reconstruction algorithm approximates the variation of the solution within each cell with a profile instead of using a constant value. The Riemann solver handles discontinuous solutions in shocks. For this experiment, we keep the timestep of the solver constant to ensure that the dynamic time-stepping algorithm does not compensate for inaccuracies resulting from truncation and thus skews our results.
@@
After we truncate the hydrodynamics, RAPTOR flags a number of operations in the reconstruction algorithm. This can be interpreted in two ways. Either, (a) the reconstruction algorithm does not respond well to truncation and should therefore be run at full precision, or (b) another part of the code is responsible for introducing small floating-point errors that only get flagged when they are amplified in the reconstruction algorithm, for example due to multiplication or cancellation. If we exclude reconstruction from truncation (but still truncate the remaining hydrodynamics code), then the errors for both density and x-velocity decrease by a small amount. Next, a number of new operations get flagged in the Riemann solver and a single operation gets flagged in the function responsible for updating the solution of the simulation. Adding the Riemann solver to the list of excluded functions drastically worsens errors, while adding the update function leaves the errors essentially unchanged.
These results indicate that no specific part of the Spark solver is responsible for numerical sensitivity in the Sedov workload and it is difficult to decide in advance whether a code section should be truncated or not. Instead, an effective truncation strategy will dynamically turn on and off based on the local smoothness of the solution in the physical regime. This small example reaffirms the need for a tool, such as RAPTOR, that can help users to understand and reason about the parts of a code that may contribute the most to floating-point errors.
@@ 7. Discussion and Outlook
We used RAPTOR to test our hypotheses about how to truncate a number of non-linear multi-physics workloads while keeping errors in check. For Sedov, we found that restricting truncation to coarse blocks worked well to keep errors low. Meanwhile for Sod, the same strategy resulted in much higher errors. These two results match Hypothesis 1. However, our Hypothesis 2 was falsified when we tried to truncate the EOS module in Cellular and found that it only converges when we run it with double precision. Finally, we evaluated the effect of precision on the evolution of the Bubble workload and found that position and shape of the split bubbles depends on the strategy and level of truncation, as the scientific intuition anticipated in our Hypothesis 3.
@@ 7.1. Overhead
[…]
@@ 7.2. Use in Hardware Co-design
We want to also highlight other adjunct usecases for RAPTOR, e.g., as part of hardware co-design (w.r.t. floating-point capabilities).
@@
We assume a simple model for a hypothetical processor with a fixed portion of chip area dedicated to floating-point processing. Given that our investigated workloads are all double precision with some portion truncated to a lower precision, we can assume that our CPU contains FPUs for double precision and one lower precision. For simplicity, we assume that the areas dedicated to each unit remain the same.
Table 4. Performance density of FPUs for various precisions (data from FPNew (Mach et al., 2021))
@@
We estimate how much processing power in floating-point operations per second (FLOP/s) a unit area of chip footprint provides for a given precision [^21] For simplicity, we assume that floating point units with different precisions are completely independent. In reality, the units often share some components.. We get an estimate for the performance density of an FPU using data from an open-source RISCV FPU implementation (i.e., FPNew (Mach et al., 2021)), which is summarized in Table 4. We extrapolate these values to get a performance density estimate for FPUs of any given precision.
Our theoretical CPU has a separate FPU for the two supported precisions (called $M_{\mathrm{dbl}}$ and $M_{\mathrm{low}}$), which, respectively, occupy an area of $A_{\mathrm{dbl}}$ and $A_{\mathrm{low}}$ and have a performance density of $P_{\mathrm{dbl}}$ and $P_{\mathrm{low}}$ operations per second per unit area (obtained above using extrapolation). Therefore, the processing capability for a specific precision is bound by a peak of $A_{i}\times P_{i}$ operations per unit time. To estimate the area that each of the two FPUs will occupy, we assume a typical CPU configuration with double and single precision capabilities in a ratio of $1:2$ (e.g. Fugaku’s A64FX (Fujitsu, [n. d.])). Using the estimated performance densities $P_{i}$, and the ratio of the compute power, we get the ratio for the areas $A_{\mathrm{dbl}}:A_{\mathrm{low}}=1.39$.
@@
To get a prediction for whether a computation is compute- or memory-bound, we build a roofline model of our processor, assuming a bandwidth of $1024$ GB/s (e.g., Fugaku’s (Fujitsu, [n. d.])).
Figure 8. Estimated speedup of Sod (cf. §4.2 and Figure 7(b)) for different truncation strategies, according to our hardware model in compute-bound and memory-bound scenarios. Results are plotted with lines (not points) for clarity. Black vertical lines mark half, float, and double precision mantissa sizes with 10, 23, and 52 bits, respectively. Irregularities around the small mantissas are due to AMR changing the ratios of truncated and non-truncated operations.:
@@ Results
Using the above model and the data collected from our experiments described in Section 6.1, we estimate speedups for different configurations of Sod and present the results in Figure 8. This is a computation-heavy application with a high operational intensity, so our roofline model predicts a compute-bound scenario. Predicted speedups for full truncation are $3.7\times$ for half precision and $2.2\times$ for single precision ($2.2\times$, $1.6\times$ memory-bound). The predicted speedups for M-1 and M-2 are lower than for M-0 because a smaller portion of the code is truncated. Consequently, speedups from low-precision operations impact the overall performance less. Irregularities around the small mantissas are caused by the adaptive behavior of AMR, which changes the ratio of truncated and non-truncated operations. This is reflected in the bar plots of Figure 7(b). For M-1, extra operations caused by AMR outweigh the speedup of the low-precision operations, resulting in net slowdowns for 4 and 5 bit mantissas. With insight from domain scientists on acceptable levels of error obtained from Figure 7(b), this information can be used to obtain an optimal FPU configuration for hardware co-design. Given that RAPTOR simplifies the collection of such information, collaborating with scientists for gathering data on the numerical behavior of software can become a powerful way to enable supercomputing centers to make informed decisions about future procurements appropriate for their workloads.
@@ 7.3. Current Limitations
For our proof-of-concept, we omitted a few quality-of-life features. For example, not all elementary functions are implemented, but adding additional functions is trivial if MPFR already supports them. RAPTOR does not truncate function calls to external, pre-compiled libraries [^22] This can be mitigated by making the function definitions available, either by providing source code or static libraries compiled with link-time optimization (LTO). or calls to dynamic function pointers. We have not tested thread-safety in general, original floating-point operands have to be FP16, FP32, or FP64, and we do not support handwritten vector assembly. Limitations outside of supporting handwritten assembly can be overcome with small engineering effort. As shown in Figure 3, the user has to manually change a few lines of code to use scoped truncation. To improve the user experience, RAPTOR can be enhanced for easier region annotation (e.g. #pragma trunc...) and to support function filtering using a configuration file (similar to profilers). Programming languages other than C/C++ and Fortran could be supported if they lower to LLVM IR. Lastly, our truncation parameters are compile-time constants, but deciding the truncation level at runtime can be achieved by compiling multiple function pointers for different truncations and conditionally using them.
@@ 8. Conclusion
This paper introduces a novel numerical profiling approach and accompanying tool, called RAPTOR, which overcomes the limitations of state-of-the-art tools in this category. Using an LLVM compiler pass and MPFR, we are able to transparently change the precision of floating-point operations in selected code regions. We create hypotheses for multiple, different workloads in the Flash-X framework for multi-physics simulations. Using our RAPTOR tool, we demonstrate how we are able to analyze and confirm/falsify these hypotheses based on collected data for the non-linear solvers which are usually hard to reason about. Our hypotheses for Sod, Sedov and Bubble experiments prove to be correct. However, our intuition turns out to be incorrect for the Cellular experiment. Currently, we can apply our tool to C/C++ and Fortran codes for CPUs and GPUs, so it supports the majority of (large-scale) scientific workloads run on supercomputers. While we outline a few limitations, none of them are fundamental to our approach and we mention how each can be mitigated with additional engineering effort. RAPTOR is a step forward towards reasoning about mixed/low precision in scientific simulations, and we hope that it finds wide adoption in the scientific computing community.
Acknowledgements.
This work was supported by the Scientific Discovery through Advanced Computing (SciDAC) program via the Office of Nuclear Physics and Office of Advanced Scientific Computing Research in the Office of Science at the U.S. Department of Energy. This work was supported by Sponsor JST SPRING , Japan Grant Number Grant #JPMJSP2180 and the RIKEN Junior Research Associate Program. This work was supported by the National Science Foundation under Grant No. 2346519, and by the Alan Turing Institute, and the U.S. Department of Energy. We gratefully acknowledge Hussein Harake and the Swiss National Supercomputing Centre (CSCS) for providing access to computing resources via the RACKlette student cluster, which was instrumental to this work.